# Distributed Multisensor ISAC

### v1

@@ I Introduction and Motivation
Integrated Sensing and Communications (ISAC) is considered one of the key features of future 6G mobile communication [ref28]. Despite different interpretations, we understand ISAC as a means of radar detection and localization of passive objects (“targets”) that are not equipped with a radio tag (“not connected”). We therefore use the term “sensing” as a synonym for “radar sensing”. The targets reveal their existence and position by radio wave reflection when properly illuminated. In contrast to well-known dedicated radar systems, ISAC exploits the inherent resources of the mobile communication system on the radio access and network level. Simple proposals on the lowest level assume to share the radio access resources, e.g. base station sites and hardware, for orthogonal transmission of communication and sensing waveforms. However, there are more advanced versions for integrating communication and sensing functions. In this paper, we solely assume that the primary system is the mobile communication network, which is extended by new sensing functionalities while taking maximum advantage from the existing communication functions. Sometimes this approach is referred to as “communication centered”. In its most resource efficient operational mode, ISAC reuses the radio resources foreseen for communication purposes also for target illumination. Obviously, the concurrent use of the same limited frequency resources for two different services is most “frequency efficient”. This allows to utilize the entire and constantly growing frequency spectrum, which is allocated for communication, also for the sensing task. This way, radar sensing can utilize the very different propagation conditions at frequencies ranging from FR1 to millimeter waves (mmWave) at FR2 and, in the future, even higher. Conversely, mobile radio can take advantage of the knowledge of the propagation environment gained from radar sensing for situation aware link adaption. The tight integration of communication and sensing functions in one system offers enormous synergy potential.
@@
Optimum design and full understanding of the possibilities and trade-offs of such fully integrated communication and sensing systems is still in its infancy. From an academic point of view, it is a challenge to determine to what extent the two fields, mobile communications and radar, can benefit and learn from each other, despite or perhaps because of the different progress and developments they have made in recent decades.
@@
The advantage of a fully integrated design of ISAC systems (often also called Integrated Communication and Sensing, ICAS) is multifaceted and goes beyond frequency reuse. We strongly believe that ISAC concepts will benefit enormously from multi-user MIMO (MU MIMO) communication concepts, as positioning is an inherently cooperative task that requires the coordination of multiple spatially distributed sensor links. Therefore, the sophisticated MU MIMO access schemes that have revolutionized the mobile radio access in communications will have a similar potential for cooperative multi-link ISAC. On the physical access level, the diverse functions of the multi-user (MU) radio access network (RAN), which are subjected to international standardization by 3GPP, already can create a huge benefit for radar sensing. This includes the full range of broadcast and orthogonal multiuser access schemes in time, frequency and space, as well as signaling techniques for channel state estimation and subsequent link adaptation, predistortion, and resource allocation. Moreover, the MU MIMO paradigm offers a perspective for distributed sensor cooperation on different levels and for building an adaptive radar sensing network where the same network is used for sensing, data transport and data fusion. With the distributed computing facilities such as the Mobile Edge Cloud (MEC) we have all resources at disposal which we may need to apply Machine Learning (ML) and Artifical Intelligence (AI) for adaptive resource allocation, target parameter estimation, and scene recognition. This way, ISAC can become a ubiquitous and cognitive radar sensing network [ref9]. These network and data processing facilities can be part of the network infrastructure and controlled by a mobile network operator. On the one hand, public and private cellular networks will offer an interesting perspective for ISAC since they are controlled and administrated by Mobile Network Operators (MNOs), which now have all tools at hand to offer an administrated radar sensing service with well-defined sensing quality of service. On the other hand, the sensing functionality can be more decentralized and organized ad-hoc if the radio network is a meshed network of direct device-to-device communication of cars or drones.
@@
Concerning the term Multiple Input Multiple Output (MIMO) we have to distinguish between collocated and distributed MIMO. While collocated MIMO in radar has a very specific meaning, which includes a pair of a sparse and a smaller dense transmitter (Tx) and receiver (Rx) array at the same platform to build a synthetic (quasi monostatic) radar array. Comparably, a distributed MIMO radar, is based on transmitters and receivers that illuminate and observe (“sense”) the target from different locations that are far away from each other. This is also referred to as bistatic or multistatic radar if several Tx-Rx pairs are involved. The advantage of the distributed MIMO architecture in ISAC is the major topic of this paper. In communications, MIMO was at first understood as a combination of antenna arrays at the Base Station (BS) and at the remote User Equipment (UE). The purpose was to establish a single data link between Tx and Rx with increased data throughput (MIMO switching or beamforming) or increased robustness (diversity). Only later multi-user MIMO was considered, which comprises several distributed UEs, usually with multiple antennas on the Rx side [ref0].
@@
The main contributions of this paper are as follows: At first we discuss basic options for distributed ISAC system concepts in conjunction with respective use cases. Then we introduce the generic architecture of a distributed multi-sensor (MS) MIMO ISAC network. We develop the basic multilink access schemes and further elaborate on a reference signal reconstruction and synchronization scheme called Cooperative Passive Coherent Location (CPCL), which we have already proposed for the first time in [ref1]. We will further work out on 3D target state vector estimation in dynamic multilink scenarios and give a comprehensive overview on multiple sensor link adaptation, coordination, adaptation and precoding in multipath environments. Finally, we offer a perspective on distributed Multi-Sensor ISAC (MS ISAC) resource allocation, data fusion and tracking. Since all the principles we propose are highly dependent on multipath propagation, we will we shortly refer to propagation aspects without referring to propagation simulation or measurement in detail.
@@ II MS ISAC Architectures and Use Cases
In this introductory section, we discuss basic system architectures for distributed ISAC networks using examples in close connection with promising use cases. To this end, we selected three network paradigms that differ in terms of access to network resources, involvement of the UE and mobility of sensor nodes. We further underline and motivate the MS ISAC architecture by referencing respective use cases that were discussed in the German 6G research initiative and coordinated by the 6G Platform [ref24]. Another white paper on use cases was published by the European standardization organization ETSI [ref24a]. The use cases selected serve as examples that benefit from distributed MS ISAC systems.
In addition to applications for car and drone traffic, we will put focus on supporting public safety and protection of critical infrastructures. We made this choice because these have not been discussed much in mobile communications so far, although they are urgently needed. We believe that these use cases directly benefit from ISAC and can initiate pilot applications, which seems most important when it comes to services to be offered by mobile operators. In the following discussion, we limit ourselves to radar detection and localization. While sensing results in other physical domains such as video, infrared, LiDAR, acoustic, etc., can be used together with radio sensing the discussion of these is beyond the scope of this paper, as we are focusing on the direct reuse of the communication radio interface for sensing. In addition, we note the radio interface can also be used to receive radio signals that are actively emitted by target objects. A malicious drone, for instance can be detected by it active radio emissions but also by radar reflections if it stops transmitting. Although this can be very important for some of the use cases below, we do not take it into account here.
@@ II-A Infrastructure-only Sensing
This variant uses only the infrastructure for target illumination, reflection detection, and data processing, which offers technical and economic advantages. The sensor nodes are stationary and integrated in a well-defined environment. This includes power supply, synchronization, access to local computational resources, real-time data exchange between nodes and to the fusion center (perhaps via a fixed front haul network) and low latency, stable internet access. Since any UE is not included in the sensing loop, commercial issues with mobile communication users are not directly affected. This may also relax privacy issues, since no user data are shared.
In terms of technical aspects, ISAC was originally considered mainly as a monostatic single site concept, consisting of a base station (gNB) equipped with an antenna array, Figure 1. This has the logistical advantage of being a standalone single-station sensing system. The array is used for Direction of Arrival (DoA) estimation. Together with Time of Flight (ToF) estimates from detected target returns (target range) we can estimate the target location. However, cross range spatial target resolution degrades with distance and the underlying monostatic radio sensing geometry would require full duplex radio access or at least collocated Tx / Rx radio interfaces (quasi monostatic). But despite of several research activities towards full duplex access, it is not yet standard in mobile radio since established communication waveforms (such as Orthogonal Frequency-Division Multiplexing (OFDM)) are not well suited for full duplex. Direct Tx to Rx feed through would be a serious problem for quasi monostatic configurations and for lower frequencies an antenna array which is necessary for DoA estimation is not always available. Moreover, for higher frequencies (FR2 and beyond), analog phased arrays respectively hybrid arrays will be used which are not directly applicable for high resolution DoA estimation. Unconventional usage of spatial precoding for DoA / Direction of Departure (DoD) estimation is described in Section V of this paper.
@@
Many of the inherent problems associated with the monostatic single station architecture can be solved with the Distributed Base Station (DBS) concept depicted in Figure 2 which changes the radar geometry from monostatic towards bistatic radio access [ref3, ref12]. Each Remote Radio Unit (RRU) receives the backscattered waves transmitted from another RRU for communication purposes in its respective downlink. For sensing, this has several advantages. To begin with, it does not require a full duplex radio interface since the direct Tx  2 Rx feed through (also called self-interference) can be controlled by the distance, position and isolation of the radio nodes and, thus, can be kept small enough. However, the sensing RRUs must be silent while one of the other RRUs is transmitting. This reduces the efficiency of communication. Moreover, we must make sure that the different RRUs have overlapping radio coverage.
@@
Figure 2: Infrastucture based sensing (Distributed gNB - multistatic). Multiple RRUs connected to a cloud-controlled baseband radio pool form a distributed base station. One RRU communicates with the UE and at the same time illuminates the scenario while other RRUs receive bistatic radar returns.
@@
As for the localization paradigm, we move from single station joint distance and DoA estimation to multiple bistatic (or multistatic) range estimation (multilateration). This requires that there is sufficient distance between the transmitting and receiving nodes. With density and spatial spread of the radio nodes deployed, we can control localization performance and coverage. Figure 2 already indicates that target position estimation uses the Excess Time of Arrival (eToA), which results from a comparison of the Line of Sight (LOS) ToF vs. ToF of the path that is routed via the target (called Bistatic Target Path (BTP)), which ends up in an ellipse indicating the target position. The focal points of the ellipse are the Tx and Rx position, respectively. To obtain the target position, we need multiple (at least three) ellipses to estimate the target position, where the accuracy depends on the length of the Tx-Rx base line. Since this multilateration localization does not require antenna arrays, it is also applicable at lower frequencies where suitable arrays may not be available. In addition, it safes measurement time, since beam search is not required, but beamforming can be additionally applied for Signal-to-Noise Ratio (SNR) gain and clutter filtering. Further advantages result from the target-related spatial diversity gain. More details will be discussed below.
@@
The distributed base station concept in Figure 2 includes synchronized RRUs (also called Remote Radio Heads (RRHs)), connected to a central, software-defined and cloud-controlled baseband radio pool (CRAN) via fixed wireless or fiber-optic fronthaul links. There are several reasons for this architecture, e.g. flexibility and easy reconfigurability, better spatial coverage, increased capacity for long-range access [ref21], [ref22], and enhanced positioning services [ref20]. For ISAC a modification of the distributed RRU concept was proposed in [ref19]. It is based on RRUs dedicated for receiving only. These “sniffer” nodes were originally proposed to submit another, well separated radio interface for quasi monostatic sensing (replacing full duplex radio interface). However, if the sniffer is deployed at a bigger distance from the RRU (e.g. several hundreds of meters away or more in a free space cellular environment), we end up with a clearly bistatic architecture or better to say a multistatic architecture if the sniffer receives also Downlink (DL) signals from other RRUs or if there are multiple sniffers deployed. This distributed, modular ISAC architecture is leveraged by the Open RAN concept that supports use of multivendor radio modules and increases functional flexibility (even out of standard) based on software-centric architectures [ref23]. Dedicated computational facilities (e.g. based on high-performance graphical processing units) for most computationally demanding radar data processing, data fusion and resource allocation fall in the scope of this concept. Also, Non-Public Networks (NPN) that are installed in wide area industrial campuses are promising for distributed infrastructure centric ISAC ecosystems.
@@
Infrastructure-only sensing seems to be a promising use case if there is a well-defined (perhaps spacious) region that needs ubiquitous and around-the-clock surveillance even if no UE is around. This can be extended outdoor industrial areas like power plants, logistic centers, railroad marshaling yards, harbors, etc., which need to be protected against intruders, spying, and terror attacks. While some high sensitive premises are already well protected by special security systems, others are not, mostly because of economic reasons. Some elements of the critical infrastructure are located in remote areas, just not observed and weakly secured. Examples are base station towers, water supply stations and energy infrastructure facilities. The latter is becoming increasingly important with the more decent power networks (power supply nodes or other de-central facilities, wind turbines, fields of photovoltaic collectors, etc.). Other examples are linear spaciously extended elements of infrastructure, such as power lines, waterways, or railway tracks. Also road traffic scenarios need more attention. While car manufacturers follow the paradigm of the so-called “ego-view” which is related to the perception of the sensors installed in the respective cars to support autonomous driving, the “regulatory view” would deserve more attention. The latter relies on a comprehensive perception of the large-scale traffic situations at traffic junctions or major intersections, traffic hot spots or other critical positions for traffic supervision by regulation authorities. Since the regulatory view cannot be generated from aggregated ego-data of cars, an independent sensor network would be necessary. The regulatory view would include recognizing dangerous situations that could lead to accidents and predicting traffic jams or blockages that could be avoided by diverting traffic flows. We can also detect illegal activities such as illegal car racing or other traffic offenses. We can detect wrong-way drivers on motorway slip roads. Another specific application is securing highway parking lots against organized cargo theft, which causes immense economic damage on German autobahn.
@@
However, fraudulent or otherwise harmful intruders can also be Unmanned Aerial Vehicless (UAVs) respectively drones. The problem is well known from air traffic disruptions at public airports by hobby drones. The problem will become much bigger with the increase of the commercial drone traffic. MNOs have already started to support Beyond Line Of Sight (BLOS) drone connectivity and offer Uncrewed Aircraft Systems Traffic Management (UTM) systems. A special airspace has been established called “U-Space”, in which rules for safe traffic are defined (see e.g. EU Drone policy 2.0). However, these UTM systems do not enough consider the chance of a violation of the UTM rules, may it be intentionally or unintentionally. Therefore, besides of cooperative traffic monitoring we would need an independent surveillance system to detect exceptional and threatening situations (“trust but verify”). However, the existing air traffic monitoring systems cannot be used for U-space surveillance because they are neither technically designed for this purpose nor can they be operated on an area-wide basis in order to economically secure the U-space level. We would need a scaled-down system being available to public and commercial drone operators everywhere. U-space surveillance is also necessary to public events against terror attacks carried out by drones and to avoid illegal observation of private and public premises by camera drones. An infrastructure-only MS ISAC system will have the technical and economic surveillance capability with ubiquitous availability. The advantages of ISAC follow from the fact that the independent surveillance capability is enabled by the same mobile communication system that also hosts the cooperative UTM system with minor infrastructure extensions. In addition, the superior range of dedicated radar is compensated for by a denser network of sensors and illuminators - another benefit of network densification.
@@
We also do not want to conceal the fact that a comprehensive solution to the problem of U-Space surveillance would require additional sensor functions. For example, radio communications to or from the drones should be monitored. An active communication link is most easy to detect and already reveals a lot about the target‘s mission. Furthermore, no transmit resources are required for target illumination. On the other hand, radar sensing has the advantage that the drone can be detected even when there is no active telemetry or data connection. It therefore makes sense to combine both methods of radio sensing, namely location of actively transmitting targets and radar location of passive targets. Finally, receiver interfaces of mobile networks can be enabled for the detection and localization of transmitters. This may be seen as another advantage of integrating radio sensing functionalities into the mobile communication network. However, a more detailed discussion of the field of radio surveillance which includes identification, recognition, localization and tracking of active radio emitters is far beyond the scope of this paper.

### v2

@@ I Introduction and Motivation
Integrated Sensing and Communications (ISAC) is considered one of the key features of future 6G mobile communication [ref28, A1, A2, A3]. Despite different interpretations, we understand ISAC as a means of radar detection and localization of passive objects (“targets”) that are not equipped with a radio tag (“not connected”). We therefore use the term “sensing” as a synonym for “radar sensing”. The targets reveal their existence and position by radio wave reflection when properly illuminated. In contrast to well-known dedicated radar systems, ISAC exploits the inherent resources of the mobile communication system on the radio access and network level. Simple proposals assume to share the radio access resources, e.g. base station sites and hardware, for orthogonal transmission of communication and sensing waveforms. However, there are more advanced versions for integrating communication and sensing functions. In this paper, we solely assume that the primary system is the mobile communication network, which is extended by new sensing functionalities while taking maximum advantage from the existing communication functions. Sometimes this approach is referred to as “communication centered”. In its most resource efficient operational mode, ISAC reuses the radio resources foreseen for communication purposes also for target illumination. Obviously, the concurrent use of the same limited frequency resources for two different services is most “bandwidth efficient”. This allows to utilize the entire and constantly growing frequency spectrum, which is allocated for communication, also for the sensing task. This way, radar sensing can utilize the very different propagation conditions at frequencies ranging from FR1 to millimeter waves (mmWave) at FR2 and, in the future, even higher. Conversely, mobile radio can take advantage of the knowledge of the propagation environment gained from radar sensing for situation aware link adaption. Summarizing, the tight integration of communication and sensing functions in one system offers enormous synergy potential.
@@
Optimum design and full understanding of the possibilities and trade-offs of such fully integrated communication and sensing systems is still in its infancy. From an academic point of view, it is a challenge to determine to what extent the two communities for mobile communications and radar can benefit and learn from each other, despite or perhaps because of the different progress and developments they have made in recent decades.
@@
The advantage of a fully integrated design of ISAC systems (often also called Integrated Communication and Sensing, ICAS) is multifaceted and goes far beyond frequency reuse. We strongly believe that ISAC concepts will benefit enormously from multi-user MIMO (MU MIMO) communication concepts, as positioning is an inherently cooperative task that requires the coordination of multiple spatially distributed sensor links. Therefore, the sophisticated MU MIMO access schemes that have revolutionized the mobile radio access in communications have a similar potential for cooperative multi-link ISAC. The idea of a distributed multi-sensor (MS) Multiple Input Multiple Output (MIMO) ISAC network which can be considered equivalent to MU MIMO and was first introduced in [EUMW, ref12, ref3]. On the physical access level, the diverse functions of the multi-user (MU) Radio Access Network (RAN), which are subjected to international standardization by 3GPP, already can create benefits for radar sensing. This includes the full range of broadcast and orthogonal multiuser access schemes in time, frequency and space, as well as signaling techniques for channel state estimation and subsequent link adaptation, predistortion, and resource allocation. Moreover, the MU MIMO paradigm offers a perspective for distributed sensor cooperation between sensors at different levels, ranging from radio access coordination to the establishment of an adaptive radar sensing network in which the same network is used for target detection, data transport, and data fusion. Further, with distributed computing facilities such as the Mobile Edge Cloud (MEC) we have all resources at our disposal which we may need to apply Machine Learning (ML) and Artifical Intelligence (AI) for adaptive resource allocation, target parameter estimation, and scene recognition. This way, ISAC can become a ubiquitous and cognitive radar sensing network [ref9]. Public and private cellular networks will offer an interesting perspective for ISAC since they are controlled and administrated by Mobile Network Operators (MNOs), which now have all tools at hand to offer an administrated radar sensing service with well-defined sensing quality of service. On the other hand, the sensing functionality can be more decentralized and organized ad-hoc if the radio network is a meshed network of direct device-to-device communication of cars or drones.
@@
Concerning the term MIMO we have to distinguish between collocated and distributed MIMO. While collocated MIMO in radar has a very specific meaning. It includes a pair of a sparse and a smaller dense transmitter (Tx) and receiver (Rx) array at the same platform to build a synthetic (quasi monostatic) radar array. Comparably, a distributed MIMO radar, is based on transmitters and receivers that illuminate and observe (“sense”) the target from distinct locations that are far away from each other. This is also referred to as bistatic or multistatic radar if several Tx-Rx pairs are involved. The advantage of the distributed MIMO architecture in ISAC is the main topic of this paper. In communications, MIMO was at first understood as a combination of antenna arrays at the Base Station (BS) and at the remote User Equipment (UE). The purpose was to establish a single data link between Tx and Rx with increased data throughput (MIMO switching or beamforming) or increased robustness (diversity). Only later multi-user MIMO was considered, which comprises several distributed UEs, usually with multiple antennas on the Rx side [ref0].
@@
The main contributions of this paper are as follows: At first we discuss basic options for distributed ISAC system concepts in conjunction with respective use cases. Then we introduce the generic architecture of a distributed MS MIMO ISAC network. We develop the basic multilink sensor access schemes and further elaborate on a reference signal reconstruction and synchronization scheme called Cooperative Passive Coherent Location (CPCL), which we have already proposed for the first time in [ref1]. We will further address the 3D target state vector estimation in dynamic multilink scenarios and give a comprehensive overview on multiple sensor link adaptation, coordination, adaptation and precoding in multipath environments. Finally, we offer a perspective on distributed Multi-Sensor ISAC (MS-ISAC) resource allocation, data fusion and tracking. Since all the principles we propose are highly dependent on multipath propagation, we will we shortly refer to propagation aspects without referring to propagation simulation or measurement in detail.
@@ II MS ISAC Architectures and Use Cases
In this introductory section, we discuss basic system architectures for distributed ISAC networks using examples in close connection with promising use cases. To this end, we selected three network paradigms that differ in terms of access to network resources, involvement of the UE and mobility of sensor nodes. We further underline and motivate the MS-ISAC architecture by referencing respective use cases that were discussed in the German 6G research initiative and coordinated by the 6G Platform [ref24]. Another white paper on use cases was published by the European standardization organization ETSI [ref24a]. The use cases selected serve as examples that benefit from distributed MS-ISAC systems.
In addition to applications for car and drone traffic, we focus on supporting public safety and protection of critical infrastructures, because these have not been discussed much in mobile communications so far, although they are urgently needed. We believe that these use cases directly benefit from ISAC and can initiate pilot applications, which is crucial when it comes to new services to be offered by mobile operators. In the following discussion, we limit ourselves to radar detection and localization. While sensor data in other physical domains such as video, infrared, LiDAR, acoustic, etc., can be used together with radio sensing the discussion of these is beyond the scope of this paper, as we are focusing on the direct reuse of the communication radio interface for sensing. In addition, we note the radio interface can also be used to receive radio signals that are actively emitted by target objects. A malicious drone, for instance can be detected by its active radio emissions but as well by radar reflections even if it stops transmitting. Although this can be very important for some of the use cases below, we do not take it into account here.
@@ II-A Infrastructure-only Sensing
This variant uses only the infrastructure for target illumination, reflection detection, and data processing, which offers technical and economic advantages. The sensor nodes are stationary and integrated in a well-defined environment. This includes power supply, synchronization, access to local computational resources, real-time data exchange between nodes and to the fusion center (perhaps via a fixed front haul network) and low latency, stable internet access. Since no UE is included in the sensing loop, commercial issues with mobile communication users are not directly affected. This may also relax privacy issues, since no user data are shared to external parties/devices.
In terms of technical aspects, ISAC was originally considered mainly as a monostatic single site concept, consisting of a base station (gNB) equipped with an antenna array, Figure 1. This has the logistical advantage of being a standalone single-station sensing system, where the array is used for Direction of Arrival (DoA) estimation. Together with Time of Flight (ToF) estimates from detected target returns (target range) one can estimate the target location. However, cross range spatial target resolution degrades with distance and the underlying monostatic radio sensing geometry would require full duplex radio access or at least collocated Tx / Rx radio interfaces (quasi monostatic). But despite of several research activities towards full duplex access, it is not yet standard in mobile radio since established communication waveforms (such as Orthogonal Frequency-Division Multiplexing (OFDM)) are not well suited for full duplex. Direct Tx to Rx feed through would be a serious problem for quasi-monostatic configurations and for lower frequencies an antenna array suitable for DoA estimation is not always available. Moreover, for higher frequencies (FR2 and beyond), analog phased arrays, or respectively hybrid arrays, will be used which are not directly applicable for high resolution DoA estimation. A related and unconventional usage of spatial precoding for DoA / Direction of Departure (DoD) estimation is described in Section V of this paper.
@@
Many of the inherent problems associated with the monostatic single station architecture can be solved with the Distributed Base Station (DBS) concept depicted in Figure 2, which changes the radar geometry from monostatic towards bistatic radio access [EUMW, ref3, ref12]. Each Remote Radio Unit (RRU) receives the backscattered waves transmitted from another RRU for communication purposes in its respective downlink. For sensing, this has several advantages. To begin with, it does not require a full duplex radio interface since the direct Tx  2 Rx feed through (also called self-interference) can be controlled by the distance, position and isolation of the radio nodes and, thus, can be kept small enough. However, the sensing RRUs must be silent while one of the other RRUs is transmitting. This reduces the efficiency of communication. Moreover, we must make sure that the different RRUs have overlapping radio coverage.
@@
Figure 2: Infrastucture based sensing (Distributed gNB- multistatic). Multiple RRUs connected to a cloud-controlled baseband radio pool form a distributed base station. One RRU communicates with the UE and at the same time illuminates the scenario, while other RRUs receive bistatic radar returns.
@@
As for the localization paradigm, we move from single station joint distance and DoA estimation to multiple bistatic (or multistatic) range estimation. Moreover, joint range/Doppler estimation for position/speed-related measurements is considered throughout this paper. Multilateration requires that there is sufficient distance between the transmitting and receiving nodes. With the density and spatial spread of the radio nodes deployed, we can control localization performance and coverage. Figure 2 already indicates that target position estimation uses the Excess Time of Flight (eToF), which results from a comparison of the Line of Sight (LOS) ToF vs. ToF of the path that is routed via the target (called Bistatic Target Path (BTP)), which ends up in an ellipse indicating the target position. The focal points of the ellipse are the Tx and Rx locations, respectively. To obtain the target position, we need multiple (at least three) intersecting ellipses, where the accuracy depends on the length of the Tx-Rx base line. Since this multilateration localization does not require antenna arrays, it is also applicable at lower frequencies where suitable arrays may not be available. In addition, it saves measurement time, since beam search is not required, but beamforming can be additionally applied for Signal-to-Noise Ratio (SNR) gain and clutter filtering. Further advantages of spatially distributed sensing result from the target-related spatial diversity gain. More details will be discussed section IV-B.
@@
The distributed base station concept in Figure 2 includes synchronized RRUs (also called Remote Radio Heads (RRHs)), connected to a central, software-defined and cloud-controlled baseband radio pool (CRAN) via fixed wireless or fiber-optic fronthaul links. There are several reasons for this architecture, e.g. flexibility and easy reconfigurability, better spatial coverage, increased capacity for long-range access [ref21, ref22], and enhanced positioning services [ref20]. For ISAC, a modification of the distributed RRU concept was proposed in [ref19]. It is based on RRUs dedicated for receiving only. These “sniffer” nodes were originally proposed to submit another, well-separated radio interface for quasi-monostatic sensing (replacing full duplex radio interface). However, if the sniffer is deployed at a bigger distance from the RRU (e.g. several hundreds of meters away or more in a free space cellular environment), we end up with a clearly bistatic architecture or better to say a multistatic architecture if the sniffer receives also Downlink (DL) signals from other RRUs or if there are multiple sniffers deployed. The proprietary link to the CRAN makes the sensor part of the infrastructure. The distributed, modular ISAC architecture is further leveraged by the Open RAN concept that supports use of multivendor radio modules and increases functional flexibility (even out of standard) based on software-centric architectures [ref23]. Dedicated computational facilities (e.g. based on high-performance graphical processing units) for most computationally demanding radar data processing, data fusion and resource allocation fall in the scope of this concept, as well as Non-Public Networks (NPN) that are installed in wide area industrial campuses are promising for distributed infrastructure-centric ISAC ecosystems.
@@
Infrastructure-only sensing seems to be a promising use case if there is a well-defined (perhaps spacious) region that needs ubiquitous and around-the-clock surveillance even if no UE is around. This can be large outdoor industrial areas like power plants, logistic centers, railroad marshaling yards, harbors, etc., which need to be protected against intruders, spying, and terror attacks. While some high sensitive premises are already well protected by special security systems, others are not, mostly because of economic reasons. Some elements of the critical infrastructure are located in remote areas, just not observed and weakly secured. Examples are base station towers, water supply stations and energy infrastructure facilities. The latter is becoming increasingly important with the more decentralized power networks (power supply nodes or other de-central facilities, wind turbines, fields of photovoltaic collectors, etc.). Other examples are linear spaciously extended elements of infrastructure, such as power lines, waterways, or railway tracks. Also road traffic scenarios need more attention. While car manufacturers follow the paradigm of the so-called “ego-view” which is related to the perception of the sensors installed in the respective cars to support autonomous driving, the “regulatory view” would deserve more attention. The latter relies on a comprehensive perception of the large-scale road situations at junctions or major intersections, accident hot spots or other critical positions for traffic supervision by regulation authorities. Since the regulatory view cannot be generated from aggregated ego-views of cars, an independent sensor network would be necessary. The regulatory view would include recognizing dangerous situations that could lead to accidents and predicting traffic jams or blockages that could be avoided by diverting traffic flows. We can also detect illegal activities such as illegal car racing or other traffic offenses as well as wrong-way drivers on motorway slip roads. Another specific application is securing highway parking lots against organized cargo theft, which causes immense economic damage on German autobahns.
@@
However, fraudulent or otherwise harmful intruders can also be Unmanned Aerial Vehicless (UAVs) respectively drones. The problem is well known from air traffic disruptions at public airports by hobby drones and will become much bigger with the increase of the commercial drone traffic. MNOs have already started to support Beyond Line Of Sight (BLOS) drone connectivity and offer Uncrewed Aircraft Systems Traffic Management (UTM) systems. A special airspace has been established called “U-Space”, in which rules for safe traffic are defined (see e.g. EU Drone policy 2.0). However, these UTM systems do not sufficiently consider the chance of a violation of the UTM rules, may it be intentionally or unintentionally. Therefore, besides of cooperative traffic monitoring we would need an independent surveillance system to detect exceptional and threatening situations (“trust but verify”). However, the existing air traffic monitoring systems cannot be used for U-space surveillance because they are neither technically designed for this purpose nor can they be operated on an area-wide basis in order to economically secure the U-space level. We would need a scaled-down system being available to public and commercial drone operators everywhere. U-space surveillance is also necessary for public events against terror attacks carried out by drones and to avoid illegal observation of private and public premises by camera drones. An infrastructure-only MS-ISAC system will have the technical and economic surveillance capability with ubiquitous availability. The advantages of ISAC follow from the fact that the independent surveillance capability is enabled by the same mobile communication system that also hosts the cooperative UTM system with minor infrastructure extensions. In addition, the superior range of dedicated radar is compensated for by a denser network of sensors and illuminators—another benefit of network densification.
@@
We also do not want to conceal the fact that a comprehensive solution to the problem of U-Space surveillance would require additional sensor functions. For example, radio communications to or from the drones should be monitored. An active communication link is most easy to detect and already reveals a lot about a target’s mission. Furthermore, no transmit resources are required for target illumination. On the other hand, radar sensing has the advantage that the drone can be detected even when there is no active telemetry or data connection. It therefore makes sense to combine both methods of radio sensing, namely localization of actively transmitting targets and radar localization of passive targets. Finally, receiver interfaces of mobile networks can be enabled for the detection and localization of transmitters. This is as another advantage of integrating radio sensing functionalities into the mobile communication network. However, a more detailed discussion of the field of radio surveillance which includes identification, recognition, localization and tracking of active radio emitters is far beyond the scope of this paper.